OpenStudy (anonymous):

The roots of x² + 26x + c = 0 are "r" and "s". Is 19r + 94s = 1994, what is the value of c?

OpenStudy (anonymous):

r+s=-26 and 19r+94s=1994. Solve for r and s. r*s=c